# Game-based Platforms for Artificial Intelligence Research

### v3

@@ Abstract
Games have been the perfect test-beds for artificial intelligence research for the characteristics that widely exist in real-world scenarios. Learning and optimisation, decision making in dynamic and uncertain environments, game theory, planning and scheduling, design and education are common research areas shared between games and real-world problems. Numerous open-source games or game-based environments have been implemented for studying artificial intelligence. In addition to single- or multi-player, collaborative or adversarial games, there has also been growing interest in implementing platforms for creative design in recent years. Those platforms provide ideal benchmarks for exploring and comparing artificial intelligence ideas and techniques. This paper reviews the game-based platforms for artificial intelligence research, provides guidance on matching particular types of artificial intelligence with suitable games for testing and matching particular needs in games with suitable artificial intelligence techniques, discusses the research trend induced by the evolution of those platforms, and gives an outlook.
Impact Statement:
Games have been playing an essential role in the development of AI techniques and education by serving as valuable test-beds. This paper provides a comprehensive review of recent games and game-based platforms for both game playing and design. The platforms discussed in this paper are comprehensively categorised based on factors such as game type, aim, number of agents, observability, programming languages, and the presence of competitions. Furthermore, the paper investigates the challenges posed by games, existing issues with game-based platforms, and the efforts in leveraging natural languages for game research. By offering a detailed roadmap and research insights, this paper serves as a valuable resource for students and researchers, facilitating their exploration of this field. Besides, the paper presents the evolution of games along with the advancement of AI techniques, thereby igniting inspiration for future investigations. Readers from industry will also discover how AI can impact games and how games can impact AI research through this survey.
Index Terms:
Game AI, platform, reinforcement learning, procedural content generation, competition
I Introduction
@@
Games have played an essential role in the rapid development of artificial intelligence (AI). As a controllable, cost-effective, and realistic simulator [1], games provide a virtual world for AI to explore. A variety of real-life problems, such as real-time decision making, scheduling, design and education, can be reflected in such virtual worlds. Games have naturally become the perfect test-beds to verify and explore the ability, generality, robustness and safety of AI. Game-based platforms have provided an ideal playground for AI researchers to study, explore, evaluate and experiment with different ideas in a controllable and safe environment. Such platforms also enable us to study and answer what-if questions that are hard to do safely in the real world. They can even stimulate researchers to ask new research questions.
@@
Various games have been used in game-based platforms, from single-player to multi-player, and from puzzles to real-time strategy games, with the emergence of diverse game genres and mechanisms. Such diversity produces a large number of diverse challenges for AI to solve. For instance, it has been demonstrated in various games (e.g., Chess [2], Atari [3], Poker [4], Go [5], StarCraft II [6], Dota 2 [7], and Texas hold’em [8]) that AI can perform at a level beyond that of humans. Besides playing games by AI, designing new games by AI is an emerging research area which helps exploit AI’s potential in creativity [9]. Games are rich in their content types, including visuals, audio, levels, rules, etc., making them challenging and diverse environments for innovative AI research [10, 11].
@@
In terms of creative game design, search-based methods, machine learning (ML) and deep learning (DL) techniques have been used for procedural content generation (PCG) where both AI creativity and human-AI co-creativity are explored [12, 13, 14, 15, 16]. Emergent abilities of large language models (LLMs) also open up new opportunities and paradigms of playing and designing games [17, 18, 19, 20, 21, 22]. More recently, Bruce et al. from DeepMind introduce Genie, the generative interactive environment that is capable of learning from videos to generate playable, endless 2D platformer games from an image prompt [23]. AI-generated contents can help enhance the generality of AI itself [11]. AI techniques can help to automate the design tasks to accelerate the process, generate novel and diverse contents, and personalise contents for individual players.
@@
This paper provides a bird’s-eye view of publicly available game-based platforms or games for AI research, aiming to provide a brief but comprehensive overview for researchers and students that want to either use AI for games or use games for AI. Fig. 1 presents the roadmap of games and platforms. Links to the source code of the reviewed games and platforms, corresponding research papers, and competitions are provided on the GitHub page [^1] All the links to the reviewed games, platforms and user instructions are collected at https://github.com/SUSTechGameAI/GameAIPlatforms..
@@
The paper is organised as follows. Section II explains the scope and methodology of this survey. Taxonomies of playing and designing tasks involving games for AI research are provided in Sections III and IV, respectively. The research trends and challenges induced by the emergence and evolution of those game-based platforms are discussed in Section V. Finally, Section VI concludes.
II Methodology of Survey
Our survey considers the interpretation of games as Liu et al. [15], “any games a human would conceivably play, including board games, card games, and any type of video games, such as arcade games, role-playing games, first-person shooters, puzzle games, and many others”. Simulators for autonomous driving or robotic control are not covered since they are specifically designed for simulating and studying robotic systems with accurate physics simulations.
@@ II Methodology of Survey
To simplify the language, from now on, we use the term “game” and “platform” to refer to a game simulator and an environment with multiple games for AI research, respectively. Yannakakis and Togelius [10] have enumerated and discussed some game-based platforms for game playing and design till 2018. Shao et al. [24] have reviewed some video game platforms for researching deep reinforcement learning till 2019. The work of Swiechowsk [1], Giannakos et al. [25], Duan et al. [26] have also touched on this topic. However, to our best knowledge, no existing paper has provided a complete review or taxonomy on recent game-based platforms, while more than 56 new ones have been released since 2018.
@@
The following survey methodology is applied. First, we have gone through the competitions organised at related conferences in AI or game AI, including NeurIPS/NIPS, IJCAI, AAAI, ICML, ECML, ACML, AIIDE, IEEE CIG, IEEE CoG, ACM FDG, ACG, GECCO, WCCI/CEC, between 2016 and February 2024, and collected corresponding platforms. The year 2016 is chosen as AlphaGo [27] has boosted the research interest in game AI. Then, we searched with Google Scholar and Web of Science using the following search terms, “game” AND (“artificial intelligence” OR “AI”) AND (“design" OR “generation”) AND (“platform” OR “interface” OR “application” OR “software”), for related papers published or available online between 2016 and February 2024, assuming that most of the earlier ones can be found in Yannakakis and Togelius [10], Shao et al. [24], and then investigated if the papers used any game-based platform for AI research. Finally, for all related papers that did not specify the link to the corresponding platform or code, we searched for the code on the authors’ personal pages, in GitHub, and using Google Search. Again, only publicly available ones are included in our survey. It’s worth mentioning that a platform or game may have been available online before the corresponding paper is published. Besides, although some commercial games are not open-source, they provide some APIs to access the games for research, e.g., Starcraft II [6]. Such games are also included.
III Playing Games by AI
@@
Games considered in this survey can be briefly categorised into tabletop games and video games. Tabletop games, including puzzles, board games and card games, usually have few requirements for image rendering, use simple objects such as boards, pieces and cards, and do not rely on digital devices. States of such games can be easily represented by structured data. Video games encompass a wide range of genres such as role-playing games, shooter games, racing games, puzzle games, and more. Unlike tabletop games, video games contain visual and audio contents based on digital devices. This poses a challenge of extracting high-dimensional but uninformative states and action features. Moreover, credit assignment in games may be challenging due to the game length, unobservable behaviours of players, combinational actions and sparse rewards. Tab. I, Tab. II and Fig. 3 categorise the games according to the game genre and the simulated physical space with which avatar(s) interact. Platforms equipped with multiple games are also introduced.
@@
TABLE I: 36 games for playing by AI. The column with header C indicates whether an AI competition is associated. Geometry Friends, available before 2016, is included as it’s used in recent research. “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent. , and denote whether the platform is deterministic, stochastic or both of them. $*{}^{*}$ highlights commercial game with publicly available API. “A”, “C”, and “M” denote adversarial, cooperative, competing as mixed scenarios.
@@
Tabletop Games (Section III-A) Board ColorShapeLinks A C# [28]
Blood bowl A Python [29]
@@
Card ben M Python [33]
@@
VGC A Python [35]
Tales of Tributes A C# [36]
@@
Video Games (Section III-B) Grid- based 2D SimCity Python [41]
@@
Snake AI A Java [43]
@@
Honor of King $*{}^{*}$ A Python [57]
Dota 2 $*{}^{*}$ M Python/ Lua [7]
StarCraft II $*{}^{*}$ A M C++/Python [6]
3D ViZDoom Python [58]
MineRL Python [59]
@@
 Single-player deterministic game  Single-player stochastic game Configurable    $*{}^{*}$ Commercial game with API
@@
TABLE II: 20 platforms with multiple games for playing by AI (cf. caption of Tab. I for reading instructions). Whether a game is adversarial or collaborative is not indicated as each platform contains a number of different games. GGP, available before 2016, is included as it’s widely used in recent research. “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent. Action-adventure game (AAG) involves video games such as puzzle and combat games.
@@
MiniWorld Python [81]
III-A Tabletop Games
@@
In ColorShapeLinks [28], players win by connecting pieces with the same colour or shape. The platform of ColorShapeLinks not only provide a visual front-end via the Unity game engine, but also a pure text-based one. ColorShapeLinks has been used for education and held as an AI competition, thus it can be easily accessed by individuals with different backgrounds. However, ColorShapeLinks is relatively simple for research purposes since good solutions can be quickly found. Blood bowl [29] simulates a fully-observable board game that two players control their own teams and are rewarded with touchdowns. At each turn, the player moves their avatars to block the team of the opposing player or pass the ball. The environment is compatible with Gym’s API for RL research [29]. In Blood bowl, the observation space consists of both spatial and non-spatial information and the action space differs across states, which leads to a new challenge in representation. Mjai [30] and Mjx [31] are partial-observable AI environments of riichi Mahjong. Thirteen tiles on the own hand and discarded cards are available for the player. If a special permutation of cards is formed, the player wins the game. Mjx is 100 times faster than Mjai in running time and is compatible with Gym’s API [31]. Diplomacy is a seven-person board game, considering both cooperation and competition [32]. In this game, AI agents are required to understand human players’ imitations and behaviours.
@@
Hearthstone: Heroes Of Warcraft is a popular commercial turn-based collectable card game (CCG). An open-source implementation version called Sabberstone [34] is released for playing and deck construction. Pokémon Video Game Championships (VGC) is a turn-based strategy game [35]. Players need to collect Pokémons with different abilities to build a team. During the battles, player agents decide the order of Pokémons and specify the skills to attack the opponents according to previous information. The work of Budzki et al. [36] introduces a deck-building card game, Tales of Tribute. Players build the deck with the same initial cards during the game through battles and trades, meeting some dynamics in a long-term strategy. Tales of Tribute also provides a graphic interface and supports implementing AI agents with C#. The three games are all CCG, where players build decks and fight against opponents, but they have different research purposes. Sabberstone and Tales of Tribute both focus on the decision agents, i.e., how to play cards. However, Tales of Tribute also considers building decks, while Sabberstone doesn’t. Meanwhile, VGC involves meta-game balance that adjusts attributes of cards during the battles. Different from some traditional card games, players of Hanabi [37] can only see the cards of other players. During the gameplay, the player can tell others some hints about their cards. As a fully cooperative game, players need to form fireworks that consist of the same colour cards in ascending order to win. Besides Poker [4] and Texas hold’em [8], Bridge is another challenging board game for AI because of its partial observability, high branching factor, cooperation and competition involved in the game [82]. 23 editions of the World Computer-Bridge Championship have been organised with the latest one in 2019222World Computer-Bridge Championship: https://bridgebotchampionship.com/. The game engine of the championship was not found by authors. However, there is a recently released game engine called ben [33] for AI research in Bridge.
@@
In most games, agents directly control the avatars to move or play cards, while in some others, agents take non-direct actions such as languages. The social card game, The Chef’s Hat [38] is also implemented based on Gym [68] and used for RL research. The winning goal of the game is to discard all hand cards [38]. AI wolf [39] is a multi-agent communication game that aims to find the wolf by talking with players. The coordination of multiple agents in partial observation only by talking is considered in the platform. Light is a text-based adventure game [40], with which dialogue-based human-AI interaction is studied.
@@
Tabletop games usually don’t require specific designs in state representation, while the action space in those games changes during an episode. For example, in Go, the state is easily constructed by a matrix, where three symbols or numbers denote white, black or empty. However, the empty space for placing stones changes in a decreasing trend, leading to a variable action space. On the other hand, tabletop games often involve multiple players with complex and even dynamic relationship (collaboration vs. competition), pose new challenges and encourage studies in various research fields such as multi-agent systems [83], game theory [84] and psychology [85].
III-B Video Games

### v4

# Games for Artificial Intelligence Research: A Review and Perspectives

@@ Abstract
Games have been the perfect test-beds for artificial intelligence research for the characteristics that widely exist in real-world scenarios. Learning and optimisation, decision making in dynamic and uncertain environments, game theory, planning and scheduling, design and education are common research areas shared between games and real-world problems. Numerous open-source games or game-based environments have been implemented for studying artificial intelligence. In addition to single- or multi-player, collaborative or adversarial games, there has also been growing interest in implementing platforms for creative design in recent years. Those platforms provide ideal benchmarks for exploring and comparing artificial intelligence ideas and techniques. This paper reviews the games and game-based platforms for artificial intelligence research, provides guidance on matching particular types of artificial intelligence with suitable games for testing and matching particular needs in games with suitable artificial intelligence techniques, discusses the research trend induced by the evolution of those games and platforms, and gives an outlook.
{IEEEImpStatement}
Games have been playing an essential role in the development of AI techniques and education by serving as valuable test-beds. This paper provides a comprehensive review of recent games and game-based platforms for both game playing and design. The platforms discussed in this paper are comprehensively categorised based on factors such as game type, aim, number of agents, observability, programming languages, and the presence of competitions. Furthermore, the paper investigates the challenges posed by games, existing issues with game-based platforms, and the efforts in leveraging natural languages for game research. By offering a detailed roadmap and research insights, this paper serves as a valuable resource for students and researchers, facilitating their exploration of this field. Besides, the paper presents the evolution of games along with the advancement of AI techniques, thereby igniting inspiration for future investigations. Readers from industry will also discover how AI can impact games and how games can impact AI research through this review.
{IEEEkeywords}
Game AI, computer games, reinforcement learning, procedural content generation, competition
1 Introduction
@@
Games have played an essential role in the rapid development of artificial intelligence (AI). As a controllable, cost-effective, and realistic simulator [1], games provide virtual worlds for AI to explore. A variety of real-life problems, such as real-time decision making [2, 3], scheduling [4], design [5, 6] and education [7], can be reflected in such virtual worlds. Games have naturally become the perfect test-beds to verify and explore the ability, generality, robustness and safety of AI [8, 9, 10]. Games or game-based platforms have provided an ideal playground for AI researchers to study, explore, evaluate and experiment with different ideas in a controllable and safe environment [5, 11]. Such games and platforms also enable us to study and answer what-if questions that are hard to do safely in the real world. They can even stimulate researchers to ask new research questions.
@@
Various games have been used in game-based research platforms, from single-player to multi-player, and from puzzles to real-time strategy (RTS) games, with the emergence of diverse game genres and mechanisms. Such diversity produces a large number of diverse challenges for AI to solve. For instance, it has been demonstrated in various games (e.g., Chess [12], Atari [2], Poker [13], Go [14], StarCraft II [15], Dota 2 [16], DouDizhu [17], Texas hold’em [18] and GuanDan [19]) that AI can perform at a level beyond that of humans. Besides playing games by AI, designing new games by AI is an emerging research area which helps exploit AI’s potential in creativity [20]. Games are rich in their content types, including visuals, audio, levels, rules, etc., making them challenging and diverse environments for innovative AI research [5, 21].
@@
In terms of creative game design, search-based methods, machine learning (ML) and deep learning (DL) techniques have been used for procedural content generation (PCG) where both AI creativity and human-AI co-creativity are explored [6, 22, 23, 24, 25, 26, 27]. Emergent abilities of large language models (LLMs) also open up new opportunities and paradigms of playing and designing games [28, 29, 30, 31, 32, 33]. More recently, Bruce et al. [34] from DeepMind introduce Genie, the generative interactive environment that is capable of learning from videos to generate playable, endless 2D platformer games from an image prompt. AI-generated contents can help to enhance the generality of AI itself [21]. AI techniques can help to automate the design tasks to accelerate the process, generate novel and diverse contents, and personalise contents for individual players [6, 26].
@@
This paper provides a bird’s-eye view of publicly available game-based platforms and games for AI research, aiming to provide a brief but comprehensive overview for researchers and students that want to either use AI for games or use games for AI. Fig. 1 presents the roadmap of games and platforms. Links to the source code of the reviewed games and platforms, corresponding research papers, and competitions are provided on the GitHub page [^1] All the links to the reviewed games, platforms and user instructions are collected at https://github.com/SUSTechGameAI/GameAIPlatforms..
@@
The paper is organised as follows. Section 2 explains the scope and methodology of this survey. Taxonomies of playing and designing tasks involving games for AI research are provided in Sections 3 and 4, respectively. The research trends and challenges induced by the emergence and evolution of those games and game-based platforms are discussed in Section 5. Finally, Section 6 concludes.
2 Methodology of Survey
@@
Our survey considers the interpretation of games as Liu et al. [6], “any games a human would conceivably play, including board games, card games, and any type of video games, such as arcade games, role-playing games, first-person shooters, puzzle games, and many others”. To simplify the language, from now on, the term “game” refers to either a game simulator or an environment with multiple games, while “game-based platforms” (or “platform” in short) specially points to the latter. Yannakakis and Togelius [5] have enumerated and discussed some games for playing and design till 2017. Shao et al. [35] have reviewed some video games for researching deep reinforcement learning (deep RL), which were released before 2019. The work of Swiechowsk [1], Giannakos et al. [36] and Duan et al. [37] have also touched on this topic. However, to our best knowledge, no existing paper has provided a complete review or taxonomy on recent games.
As a side note, the fields of game AI, autonomous driving,robotics and etc., share some similar challenges [38, 39]. On one hand, some of the games, simulators or virtual environments may be used by one or more fields. On the other hand, the AI approaches and techniques widely used in one field can be applied to another. However, to be more focused in this survey, virtual environments specifically designed for simulating and studying autonomous driving and robotic systems, such as Carla [40], Omniverse [41] and Mujoco [42] with accurate physics simulations, are not covered.
@@
The following survey methodology is applied. First, we have gone through the competitions organised at related conferences in AI or game AI, including NeurIPS/NIPS, IJCAI, AAAI, ICML, ECML, ACML, AIIDE, IEEE CIG, IEEE CoG, ACM FDG, ACG, GECCO, WCCI/CEC, between 2016 and February 2024, and collected corresponding games/ platforms. Then, we searched with Google Scholar and Web of Science using the following search terms, “game” AND (“artificial intelligence” OR “AI”) AND (“design" OR “generation”) AND (“platform” OR “interface” OR “application” OR “software”), for related papers published or available online between 2016 and February 2024, assuming that most of the earlier ones can be found in Yannakakis and Togelius [5], Shao et al. [35], and then investigated if the papers used any game-based platform for AI research. Finally, for all related papers that did not specify the link to the corresponding platform or code, we searched for the code on the authors’ personal pages, in GitHub, and using Google Search. The year 2016 is chosen for the following reasons. AlphaGo [43] introduced in 2016 is a milestone for AI research. Its success presents the potential of several AI approaches and techniques, such as deep RL and search methods, in mastering complex games with vast search spaces and strategic depth beyond the human level, and inspires not only research on game AI [44, 45] but also interdisciplinary research, such as AI for science [46, 47] and AI for social interaction [48]. As discussed previously, most of the related games published before 2017 have been covered by [5, 35]. However, lots of new games have emerged as shown in this survey. Again, only publicly available ones are included. It’s worth mentioning that a platform or game may have been available online before the corresponding paper is published. Besides, although some commercial games are not open-source, they provide some APIs to access the games for research, e.g., StarCraft II [15]. Such games are also included.
3 Playing Games by AI
@@
Games considered in this survey can be briefly categorised into tabletop games and video games. Tabletop games, including puzzles, board games and card games, usually have few requirements for image rendering, use simple objects such as boards, pieces and cards, and do not rely on digital devices. States of such games can be easily represented by structured data. Video games encompass a wide range of genres such as role-playing games, shooter games, racing games, puzzle games, and more. Unlike tabletop games, video games contain visual and audio contents based on digital devices. This poses a challenge of extracting high-dimensional but uninformative states and action features. Moreover, credit assignment in games may be challenging due to the game length, hidden behaviours of players, combinatorial actions and sparse rewards. Tab. 1, Tab. 2 and Fig. 3 categorise the games and platforms according to the game genre, research aims, the number of agents and observability of the simulated physical space with which agent(s) interact, as well as programming language and if they have been used by any AI-related competitions.
@@
Table 1:  40 games for playing by AI. The column with header C indicates whether an AI competition is associated. Geometry Friends, available before 2016, is included as it’s used in recent research. “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent. , and denote whether the platform is deterministic, stochastic or both. highlights commercial games with publicly available API.
@@
Tabletop Games (Section 3.1) Board ColorShapeLinks A C# [49]
@@
Card VGC A Python [54]
Tales of Tributes A C# [55]
@@
Fireplace A Python [57]
ben M Python [58]
@@
Video Games (Section 3.2) Grid- based 2D Snake AI A Java [63]
SimCity Python [64]
@@
Pommerman M Java/Python [66]
Neural MMO M Python [67]
@@
AIBIRDS Java/Python [78]
nsh Python/Lua [79]
Honor of King  A Python [80]
Dota 2  M Python/Lua [16]
Dunk City Dynasty  M Python [81]
StarCraft II  A M C++/Python [15]
3D MineRL Python [82]
@@
Google Research Football M Python [84]
ViZDoom Python [85]
@@
 Single-player deterministic game  Single-player stochastic game Configurable Commercial game with API
@@
Table 2: 20 platforms with multiple games for playing by AI (cf. caption of Tab. 1 for reading instructions). Whether a game is adversarial or collaborative is not indicated as each platform contains a number of different games. GGP, available before 2016, is included as it’s widely used in recent research. “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent. Action-adventure game (AAG) involves video games such as puzzles and combat games.
@@
3.1 Tabletop Games
@@
In ColorShapeLinks [49], players win by connecting pieces with the same colour or shape. The platform of ColorShapeLinks provide not only a visual front-end via the Unity game engine, but also a pure text-based one. ColorShapeLinks has been used for education and a testbed of an AI competition. It can be easily accessed by individuals with different backgrounds. However, ColorShapeLinks is relatively simple for research purposes since good solutions can be quickly found. Mjai [50] and Mjx [51] are partial-observable AI environments of riichi Mahjong. Thirteen tiles on the own hand and discarded cards are available for the player. If a special permutation of cards is formed, the player wins the game. Mjx is 100 times faster than Mjai in running time and is compatible with Gym’s API [51]. Blood bowl [52] simulates a fully-observable board game that two players control their own teams and are rewarded with touchdowns. At each turn, the player moves their avatars to block the team of the opposing player or pass the ball. The environment is compatible with Gym’s API for RL research [52]. In Blood bowl, the observation space consists of both spatial and non-spatial information and the action space differs across states, which leads to a new challenge in representation. Diplomacy is a seven-person board game, considering both cooperation and competition [53]. In this game, AI agents are required to understand human players’ imitations and behaviours.
@@
Pokémon Video Game Championships (VGC) is a turn-based strategy game [54]. Players need to collect Pokémons with different abilities to build a team. During the battles, player agents decide the order of Pokémons and specify the skills to attack the opponents according to previous information. The work of [55] introduces a deck-building card game, Tales of Tribute. Players build the deck with the same initial cards during the game through battles and trades, meeting some dynamics in a long-term strategy. Tales of Tribute also provides a graphic interface and supports implementing AI agents with C#. Hearthstone: Heroes Of Warcraft is a popular commercial turn-based CCG. There are open-source implementations, e.g., Sabberstone [56] and Firepalce [57], for playing. The three games are all collectable card games (CCG), where players build decks and fight against opponents, but they have different research purposes. Sabberstone, Firepalce and Tales of Tribute focus on researching decision-making agents, i.e., how to play cards. However, Tales of Tribute can also be used to study building decks by AI. Meanwhile, VGC involves meta-game balance that adjusts attributes of cards during the battles. Bridge is another challenging board game for AI because of its partial observability, high branching factor, cooperation and competition [107]. 23 editions of the World Computer-Bridge Championship have been organised with the latest one in 2019222World Computer-Bridge Championship: https://bridgebotchampionship.com/. The game engine of the championship was not found by authors. However, there is a recently released game engine called ben [58] for AI research in Bridge. Different from some traditional card games, players of Hanabi [59] can only see the cards of other players. During the gameplay, the player can tell others some hints about their cards. As a fully cooperative game, players need to form fireworks that consist of the same colour cards in ascending order to win.
@@
In most games, agents directly control avatars to move or play cards, while in some others, agents take non-direct actions such as languages. The social card game, The Chef’s Hat [60] is also implemented based on Gym [93] and used for RL research. The winning goal of the game is to discard all hand cards [60]. AI wolf [61] is a multi-agent communication game that aims to find the wolf by talking with players. The coordination of multiple agents in partial observation only by talking is considered in the platform. Light is a text-based adventure game [62], with which dialogue-based human-AI interaction is studied.
@@
Usually, no specific design is required for representing states in tabletop games, while the action space in those games changes during an episode. For example, in the game of Go, the game state is easily constructed by a matrix, where three symbols or digits denote white, black or empty. However, the empty space for placing stones changes in a decreasing trend, leading to a variable action space. On the other hand, tabletop games often involve multiple players with complex and even dynamic relationship (collaboration vs. competition), pose new challenges and encourage studies in various research fields such as multi-agent RL [108, 109, 110], game theory [111, 112] and psychology [113].
3.2 Video Games